OpenStudy (anonymous):

A dartboard has three concentric circles and the radius of the inner circle is 1. If the probability of landing in each of the three regions is the same, find the EXACT value of the radii of the other two circles. SHOW WORK!

OpenStudy (jack1):

therefore the area of each circle is the same so A of Circle = pi x r^2

OpenStudy (jack1):

area of 2nd circle = (pi x r^2) - area of circle 1

OpenStudy (jack1):

area of 3rd circle = (pi x r^2) - (area of circle 1 + area of circle 2)

OpenStudy (jack1):

as circles have equal area: Area of circle 1= area of circle 2 = area of circle 3

OpenStudy (anonymous):

What is the area?

OpenStudy (jack1):

you have the radius, you tell me

OpenStudy (anonymous):

Well pi (1)2 is 1pi so they are all 1pi?

OpenStudy (jack1):

yep, perfect

OpenStudy (anonymous):

Huh, that was more simple than I thought. Thank you very much.

OpenStudy (jack1):

the q was to find out the radius tho

OpenStudy (jack1):

so radius of circle 1 is given : 1

OpenStudy (jack1):

radius of circle 2 is... area of 2nd circle = (pi x r^2) - area of circle 1 3.14 = (pi x r^2) - 3.14 2 pi = pi x r^2 2 = r^2 r = sqr rt of 2

OpenStudy (jack1):

following ok so far?

OpenStudy (jack1):

so tell me what you get for radius of circle 3?

OpenStudy (anonymous):

Urm, do you subtract 3.14 from the area of the first and second circle?

OpenStudy (jack1):

start with this equation: area of 3rd circle = (pi x r^2) - (area of circle 1 + area of circle 2)

OpenStudy (jack1):

sub in 3.14 wherever you see "area of circle #"

OpenStudy (anonymous):

3pi = pir2 3pi = r^2? r= sqr of 3?

OpenStudy (jack1):

perfect!

OpenStudy (jack1):

your second line's a typo, yeah? its supposed to read 3 = r^2?

OpenStudy (anonymous):

Yeah, opps.

OpenStudy (jack1):

all good, well done

OpenStudy (anonymous):

Thank you so much for helping. That was very simple and easy :).

OpenStudy (jack1):

was a pleasure